OpenStudy (anonymous):

The ratio of the measures of two supplementary angles is 7:3. Find the measures of the angles.

OpenStudy (anonymous):

If "The ratio of the measures of two supplementary angles is 7:3", then 7x+3x = 180 Solve for x. Then use this to find 7x and 3x

OpenStudy (lncognlto):

7+3 = 10, so the angle, which equals 180, is in 10 parts. 180/10 = 18. So the angles are then 18*7 and 18*3

OpenStudy (anonymous):

A. 126° and 54° B. 72° and 18° C. 159° and 21° D. 69° and 21°

OpenStudy (lncognlto):

18*7 = 126 18*3 = 54 So then A...